# Higher-order triadic percolation on random hypergraphs

### v1

@@ V.2 Generalization to regulated nodes
Figure 15: Panel (a) provides a comparison among the phase diagram of ordinary hypergraph percolation with node deactivation, indicated here as node percolation (NP) and with hyperedge deactivation, indicated here as hyperedge percolation (HP). Panel (b) provides a comparison among the phase diagram of HOTP and HONDP. Solid lines indicate theoretical predictions, symbols indicate Monte Carlo simulations conducted on hypergraphs of $N=10^{4}$ nodes. The structural hypergraph has a Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=4$ and the hyperedges with fixed cardinality $m=5$. In panel (b) the positive regulatory interactions have a Poisson distribution with an average $c^{+}=4$ while negative regulatory interactions are absent.
@@
Secondly we compare the properties of case (a) and case (b) HONDP and we observe that when the hyperedge cardinality is constant, and is given by $m$, these two models have different responses to the increase on $m$. Both percolation processes display a discontinuous phase transition, but the critical threshold of monotonically decreasing with $m$ in case (a) where Step 1 is defined in terms of the hyperedge percolation process, while is and increasing function of $m$ in case (b) where Step 2 is defined in terms of the cooperative hyperedge percolation process (see Figure 16). This different behavior is inherently related to the different nature of the two models as the rules of cooperative hyperedge percolation are determining the increased fragility of hypergraphs with hyperedges of larger cardinality.
@@
Figure 16: Phase diagrams of HONDP in case (a) implementing simple node hyperedge percolation (panel (a)) and in case (b) implementing cooperative node hyperedge percolation (panel (b)) for hypergraphs with hyperedges of fixed cardinality $m$ and in absence of negative node regulations. The numerical predictions (solid lines) obtained from Eq. 42 and Eq. 43) are compared with Monte Carlo simulations (symbols). In both cases we observe a discontinuous hybrid transition, however in case (a) the percolation threshold decreases with increasing values of $m$ while in case (b) it increases. This demonstrates that while in case (a) HODNP the hypergraphs become more robust by increasing the value of $m$ in case (b) HODNP the hypergraphs become more fragile. The considered structural hypergraph has $N=10^{5}$ nodes and a Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=6$. The Poisson degree distribution $\hat{P}^{+}(\kappa)$ has average positive regulatory degree $c^{+}=4$ while the negative regulations are absent.
@@ VI Conclusion
In this study, we propose a comprehensive theoretical framework that merges percolation theory with nonlinear dynamics to examine hypergraphs featuring a time-varying giant component. In higher-order triadic percolation (HOTP) the dynamics is orchestrated by higher-order triadic interactions (HOTIs) that can upregulate or downregulate hyperedges turning percolation into a fully-fledged dynamical process. Specifically, the order parameter of HOTP, indicating the fraction of nodes in the giant component, undergoes, in the most general scenario, period-doubling and a route to chaos. Here we provide an in depth study of the critical behavior of this model characterizing the emergence of discontinuous transition when only positive HOTIs are allowed; the emergence of both continuous and discontinuous phase transitions connected by tricritical points when only negative HOTIs are allowed; and proving that, in presence of both positive and negative HOTIs, the order parameter of HOTP undergoes a route to chaos in the universality class of the logistic map.
Higher-order triadic interactions, however, can also be hierarchically nested giving rise to hierarchical higher-order triadic interactions (HHOTIs) that have been observed for instance in ecological networks. In this scenario, it is important to consider a relevant variation of HOTP that we call hierarchical higher-order triadic percolation (HHOTP). Due to the improved combinatorial complexity of the regulatory interactions allowed in this scenario, we observe that the critical properties of HHOTP can be significantly more complex than those of HOTP. Among the most relevant differences between the two models we mention two. First, HHOTP can display a route to chaos of the order parameter also when only negative HHOTIs are allowed. Secondly, the route to chaos associated to HHOTP is not generally in the universality class of the logistic map.
We conclude our work presenting two other generalizations of HOTP: interdependent higher-order triadic percolation (IHOTI) and higher-order node dynamic percolation (HONDP).
The underlying percolation process in IHOTI is the cooperative hypergraph percolation which already displays a discontinuous percolation transition in absence of HOTIs. This impedes the route to chaos of the order parameter for certain parameter values, where only stationary states of period-2 oscillations of the order parameter can be observed.
@@
In conclusion, this work provides a comprehensive framework for studying hypergraphs with time-varying connectivity, revealing the complex interplay between higher-order network structures and their dynamics. The results might shed light on the dynamics of climate networks, biological and brain networks.
@@ Appendix A Universality class of the route to chaos associated with HOTP
In this Appendix our aim is to determine the universality class associated to the route to chaos associated with HOTP. A classic result of chaos theory is that an unimodal continuous one-dimensional map undergoes a route to chaos in the universality class of the logistic map if it has a uniquely differentiable quadratic maximum [49, 50].
@@
This implies that the map $h_{m,p}(R)$ admits a maximum only if the equations implementing the regulatory step are not monotonic, i.e. $p_{H}^{(t)}$ is not a monotonic function of $R^{(t-1)}$. Thus this implies that both positive and negative HOTIs need to be present with a non vanishing density. In order to prove that the maximum $R=\tilde{R}$ is quadratic, we consider the second derivative of the map
@@
This equation implies that, given the inequalities (52), the sign of $h^{\prime\prime}_{m,p}(\tilde{R})$ depends only on the form of ${\partial^{2}F_{4}}/{\partial(R^{(t-1)})^{2}}$. In particular, the map $h_{m,p}(R)$ will always have a quadratic maximum . This demonstrates that the map $h_{m,p}(R)$ will always have a quadratic maximum, if $p_{H}^{(t)}=F_{4}(R^{(t-1)})$ displays a quadratic maximum. Thus in such general hypothesis the order parameter $R=R^{(t)}$ of the HOTP displays a route to chaos in the universality class of the logistic map. Note that this implies that the universality class of HOTP is independent of the hypergraph structure and that also changing the regulatory degree distribution might not affect this result as long as $p_{H}^{(t)}=F_{4}(R^{(t-1)})$ displays a quadratic maximum at a non-zero value of $R$.
In this work we focus mostly in the relevant case in which the regulatory (positive and negative) degree distributions are Poisson distributions with average degrees $c^{\pm}$ respectively. This case satifies the hypothesis for observing a route to chaos of the order parameter of HOTP in the universlaity class of the logistic map. In fact we have
@@
Figure 17: Illustration of the functions $f_{m}(p_{H})$ (solid line) and $g_{p}(R)$ (dashed lines) determining the critical properties of HOTP when hyperedges are not regulated with probability $\rho_{0}$. In panel (a) we consider values of $\rho_{0}$ above the tricritical point, where a continuous phase transition is observed. In panel (b) we consider values of $\rho_{0}$ below the tricritical point where a discontinuous phase transition is observed. In panel (c) we consider values of $\rho_{0}$ at the tricritical point. In each panels, $p_{c}$ indicates the critical threshold for discontinuous transition (panel (a)) continuous transition (panel (b)) and tricritical phase transition (panel (c)). The structural hypergraph has fixed hyperedge cardinality $m=2$ and a Poisson hyperdegree distribution with an average $c=3$. The regulatory network is formed exclusively by positive regulations with a Poisson regulatory degree distribution with an average $c^{+}=2$.
@@ Appendix B HOTP with partial regulation
In the main text we have shown that HOTP in absence of negative regulations can only display a discontinuous hybrid phase transition, while HOTP in absence of positive regulation can display either continuous transition or period-2 oscillations. In order to reveal the nature of the difference between these two scenarios, we discuss here a general scenario of HOTP with partial regulation where hyperedges are not regulated with probability $\rho_{0}$. Finally we discuss also the effect of partial regulation in HONDP.
B.0.1 HOTP in absence of negative regulations
Let us consider HOTP where hyperedges are not regulated with probability $\rho_{0}$, i.e. the HOTIs are present with a probability $1-\rho_{0}$. the probability of retaining a hyperedge $p_{H}^{(t)}$ at time $t$ follow the following generalization of Eq. 8,
@@ B.0.1 HOTP in absence of negative regulations
When $\rho_{0}=1$, we recover ordinary hypergraph percolation with random hyperedge deactivation that displays a second-order continuous phase transition. When $\rho_{0}=0$, we recover the HOTP in absence of negative regulation which displays a discontinuous hybrid phase transition.
Let us investigate the general scenario when $10<\rho_{0}<$ 1. First we observe that when $R=0$, $p_{H}=\rho_{0}p$. Let us denote the percolation threshold for for simple hyperedge percolation as $p_{H}^{c}$ (the blue line in Fig. 17) and let us adopt the map notation for the dynamics,
@@
We observe that $\rho_{0}>0$ for any $c^{+}$ and $m$. Therefore in the case of $HOTP$ in absence of negative regulations, $\rho_{0}=0$, the transition will always be discontinuous.
B.0.2 HHOTP in absence of positive regulations
@@
Assuming the hypergraph has a Poisson structuralhyperdegree distribution with an average $c$, a Poisson regulatory degree distribution with an average $c^{-}$ and a fixed hyperedge cardinality $m$,
@@
Thus, when $\rho_{0}=0$, we have $c^{-}=(m-1)/2$. Therefore, in the case of HOTP in absence of positive regulation, as long as $c^{-}<(m-1)/2$, the transition is always continuous.
@@ B.0.3 HONDP with partial regulation
We we discuss also HONP when hyperedges are not regulated with probability $\rho_{0}$ observing a critical behavior similar to the one of HOTP with partial regulation discussed in the previous paragraphs of this Appendix. The major difference between HONP and HOTP is that the map $R=f_{m}(p)$ implements hypergraph percolation (case (a)) or cooperative hypergraph percolation (case (b)) with deactivation of the nodes instead of hypergraph percolation with hyperedge deactivation. Let us discuss the case (a) as an example. In this case we have
@@
Thus for HONDP in absence of negative regulations, the tricritical point is given by
@@
For HONDP in absence of positive regulation, we have instead that the tricritical point is given by

### v2

@@ V.2 Generalization to regulated nodes
Figure 16: Panel (a) provides a comparison among the phase diagram of ordinary hypergraph percolation with node deactivation, indicated here as node percolation (NP), and with hyperedge deactivation, indicated here as hyperedge percolation (HP). Panel (b) provides a comparison among the phase diagrams of HOTP and HONDP. Solid lines indicate theoretical predictions, symbols indicate Monte Carlo simulations conducted on hypergraphs of $N=10^{4}$ nodes. The structural hypergraph has a Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=4$ and the hyperedges with fixed cardinality $m=5$. In panel (b) the positive regulatory interactions have a Poisson distribution with an average $c^{+}=4$ while negative regulatory interactions are absent.
@@
Secondly, we compare the properties of case (a) and case (b) HONDP and we observe that when the hyperedge cardinality is constant, and is given by $m$, these two models have different responses to the increase on $m$. Both percolation processes display a discontinuous phase transition, but the critical threshold of monotonically decreasing with $m$ in case (a) where Step 1 is defined in terms of the hyperedge percolation process, while is an increasing function of $m$ in case (b) where Step 2 is defined in terms of the cooperative hyperedge percolation process (see Figure 17). This different behavior is inherently related to the different nature of the two models as the rules of cooperative hyperedge percolation determine the increased fragility of hypergraphs with hyperedges of larger cardinality.
@@
Figure 17: Phase diagrams of HONDP in case (a) implementing simple node hyperedge percolation (panel (a)) and in case (b) implementing cooperative node hyperedge percolation (panel (b)) for hypergraphs with hyperedges of fixed cardinality $m$ and in the absence of negative node regulations. The numerical predictions (solid lines) obtained from Eq. (42) and Eq. (43)) are compared with Monte Carlo simulations (symbols). In both cases we observe a discontinuous hybrid transition, however in case (a) the percolation threshold decreases with increasing values of $m$ while in case (b) it increases. This demonstrates that while in case (a) HODNP the hypergraphs become more robust by increasing the value of $m$ in case (b) HODNP the hypergraphs become more fragile. The considered structural hypergraph has $N=10^{5}$ nodes and a Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=6$. The Poisson degree distribution $\hat{P}^{+}(\kappa)$ has average positive regulatory degree $c^{+}=4$ while the negative regulations are absent.
@@ VI Conclusion
In this study, we propose a comprehensive theoretical framework that merges percolation theory with nonlinear dynamics to examine hypergraphs featuring a time-varying giant component. In higher-order triadic percolation (HOTP) the dynamics is orchestrated by higher-order triadic interactions (HOTIs) that can upregulate or downregulate hyperedges turning percolation into a fully-fledged dynamical process. Specifically, the order parameter of HOTP, indicating the fraction of nodes in the giant component, undergoes, in the most general scenario, period doubling and a route to chaos. Here we provide an in-depth study of the critical behavior of this model characterizing the emergence of discontinuous transition when only positive HOTIs are allowed; the emergence of both continuous and discontinuous phase transitions connected by tricritical points when only negative HOTIs are allowed; and proving that, in the presence of both positive and negative HOTIs, the order parameter of HOTP undergoes a route to chaos in the universality class of the logistic map.
Higher-order triadic interactions, however, can also be hierarchically nested, giving rise to hierarchical higher-order triadic interactions (HHOTIs) that have been observed for instance in ecological networks. In this scenario, it is important to consider a relevant variation of HOTP that we call hierarchical higher-order triadic percolation (HHOTP). Due to the improved combinatorial complexity of the regulatory interactions allowed in this scenario, we observe that the critical properties of HHOTP can be significantly more complex than those of HOTP. Among the most relevant differences between the two models, we highlight two of them. First, HHOTP can display a route to chaos of the order parameter also when only negative HHOTIs are allowed. Secondly, the route to chaos associated with HHOTP is not generally in the universality class of the logistic map.
We conclude our work by presenting two other generalizations of HOTP: interdependent higher-order triadic percolation (IHOTI) and higher-order node dynamic percolation (HONDP).
The underlying percolation process in IHOTI is the cooperative hypergraph percolation which already displays a discontinuous percolation transition in the absence of HOTIs. This impedes the route to chaos of the order parameter for certain parameter values, where only stationary states of period-2 oscillations of the order parameter can be observed.
@@
In conclusion, this work provides a comprehensive framework for studying hypergraphs with time-varying connectivity, revealing the complex interplay between higher-order network structures and their dynamics. The results might shed light on the dynamics of climate networks, and biological and brain networks.
@@ Appendix A Universality class of the route to chaos associated with HOTP
In this Appendix our aim is to determine the universality class associated with the route to chaos associated with HOTP. A classic result of chaos theory is that an unimodal continuous one-dimensional map undergoes a route to chaos in the universality class of the logistic map if it has a uniquely differentiable quadratic maximum [51, 52].
@@
This implies that the map $h_{m,p}(R)$ admits a maximum only if the equations implementing the regulatory step are not monotonic, i.e. $p_{H}^{(t)}$ is not a monotonic function of $R^{(t-1)}$. Thus this implies that both positive and negative HOTIs need to be present with a non-vanishing density. In order to prove that the maximum $R=\tilde{R}$ is quadratic, we consider the second derivative of the map
@@
This equation implies that, given the inequalities (52), the sign of $h^{\prime\prime}_{m,p}(\tilde{R})$ depends only on the form of ${\partial^{2}F_{4}}/{\partial(R^{(t-1)})^{2}}$. In particular, the map $h_{m,p}(R)$ will always have a quadratic maximum. This demonstrates that the map $h_{m,p}(R)$ will always have a quadratic maximum, if $p_{H}^{(t)}=F_{4}(R^{(t-1)})$ displays a quadratic maximum. Thus in such general hypotheses the order parameter $R=R^{(t)}$ of the HOTP displays a route to chaos in the universality class of the logistic map. Note that this implies that the universality class of HOTP is independent of the hypergraph structure and that also changing the regulatory degree distribution might not affect this result as long as $p_{H}^{(t)}=F_{4}(R^{(t-1)})$ displays a quadratic maximum at a non-zero value of $R$.
In this work, we focus mostly on the relevant case in which the regulatory (positive and negative) degree distributions are Poisson distributions with average degrees $c^{\pm}$ respectively. This case satisfies the hypothesis for observing a route to chaos of the order parameter of HOTP in the universality class of the logistic map. In fact, we have
@@
Figure 18: Illustration of the functions $f_{m}(p_{H})$ (solid line) and $g_{p}(R)$ (dashed lines) determining the critical properties of HOTP when hyperedges are not regulated with probability $\rho_{0}$. In panel (a) we consider values of $\rho_{0}$ above the tricritical point, where a continuous phase transition is observed. In panel (b) we consider values of $\rho_{0}$ below the tricritical point where a discontinuous phase transition is observed. In panel (c) we consider values of $\rho_{0}$ at the tricritical point. In each panel, $p_{c}$ indicates the critical threshold for discontinuous transition (panel (a)) continuous transition (panel (b)), and tricritical phase transition (panel (c)). The structural hypergraph has fixed hyperedge cardinality $m=2$ and a Poisson hyperdegree distribution with an average $c=3$. The regulatory network is formed exclusively by positive regulations with a Poisson regulatory degree distribution with an average $c^{+}=2$.
@@ Appendix B HOTP with partial regulation
In the main text, we have shown that HOTP in the absence of negative regulations can only display a discontinuous hybrid phase transition, while HOTP in the absence of positive regulation can display either continuous transition or period-2 oscillations. In order to reveal the nature of the difference between these two scenarios, we discuss here a general scenario of HOTP with partial regulation where hyperedges are not regulated with probability $\rho_{0}$. Finally, we discuss also the effect of partial regulation in HONDP.
B.0.1 HOTP in the absence of negative regulations
Let us consider HOTP where hyperedges are not regulated with probability $\rho_{0}$, i.e. the HOTIs are present with a probability $1-\rho_{0}$. the probability of retaining a hyperedge $p_{H}^{(t)}$ at time $t$ follow the following generalization of Eq. (8),
@@ B.0.1 HOTP in the absence of negative regulations
When $\rho_{0}=1$, we recover ordinary hypergraph percolation with random hyperedge deactivation that displays a second-order continuous phase transition. When $\rho_{0}=0$, we recover the HOTP in the absence of negative regulation which displays a discontinuous hybrid phase transition.
Let us investigate the general scenario when $10<\rho_{0}<$ 1. First, we observe that when $R=0$, $p_{H}=\rho_{0}p$. Let us denote the percolation threshold for for simple hyperedge percolation as $p_{H}^{c}$ (the blue line in Fig. 18) and let us adopt the map notation for the dynamics,
@@
We observe that $\rho_{0}>0$ for any $c^{+}$ and $m$. Therefore in the case of $HOTP$ in the absence of negative regulations, $\rho_{0}=0$, the transition will always be discontinuous.
B.0.2 HHOTP in the absence of positive regulations
@@
Assuming the hypergraph has a Poisson structural hyperdegree distribution with an average $c$, a Poisson regulatory degree distribution with an average $c^{-}$, and a fixed hyperedge cardinality $m$,
@@
Thus, when $\rho_{0}=0$, we have $c^{-}=(m-1)/2$. Therefore, in the case of HOTP in the absence of positive regulation, as long as $c^{-}<(m-1)/2$, the transition is always continuous.
@@ B.0.3 HONDP with partial regulation
We discuss also HONP when hyperedges are not regulated with probability $\rho_{0}$ observing a critical behavior similar to the one of HOTP with partial regulation discussed in the previous paragraphs of this Appendix. The major difference between HONP and HOTP is that the map $R=f_{m}(p)$ implements hypergraph percolation (case (a)) or cooperative hypergraph percolation (case (b)) with deactivation of the nodes instead of hypergraph percolation with hyperedge deactivation. Let us discuss the case (a) as an example. In this case, we have
@@
Thus for HONDP in the absence of negative regulations, the tricritical point is given by
@@
For HONDP in the absence of positive regulation, we have instead that the tricritical point is given by